### v2

# Molecular mechanism of heterogeneous ice nucleation in the atmosphere

Abstract
Mineral dust aerosols strongly influence Earth’s climate by acting as ice-nucleating particles (INPs). Feldspar minerals, particularly K-feldspar, are recognized as dominant INPs, and a previous study attributed this behavior to (100) surfaces exposed at defects. Using machine-learning molecular dynamics simulations, we systematically investigate ice nucleation on all K-feldspar surfaces. We identify the (110) surface, exposed at defects such as steps, as the most active plane for ice formation. This surface uniquely structures interfacial water into an arrangement resembling that on the (110) surface of cubic ice, providing an optimal template for nucleation. Using advanced sampling, we directly observe the formation of clusters with cubic-ice structure and their orientation agrees with experiment. These results provide a molecular-level explanation of how ice forms in our planet’s atmosphere.
Introduction
Ice nucleation in the atmosphere has an important impact on the climate of our planet, affecting precipitation formation, the hydrological cycle, and Earth’s radiation budget  (?, ?). Understanding the mechanism of this process is essential for climate modeling and prediction, a subject of increasing relevance due to anthropogenic climate change  (?), as well as for multiple applications such as cloud seeding, the production of artificial snow, anti-icing strategies for aircraft, and frost protection in agriculture. Ice can nucleate from liquid water via two distinct pathways  (?). The first is homogeneous ice nucleation, which occurs in pure supercooled water at sufficiently low temperatures, namely, below -38°C. More commonly, ice forms via heterogeneous ice nucleation, where it nucleates on a wide variety of atmospheric aerosols known as ice-nucleating particles (INPs)  (?, ?). These particles provide sites or surfaces where ice can readily form, thereby significantly reducing the free energy barrier for the transformation (?). As a result, heterogeneous ice nucleation can occur at warmer conditions, typically between -38°C and -10°C  (?).
The importance of each type of INP was not fully understood until 2013 when landmark work  (?, ?) demonstrated that potassium (K)-feldspar dominates atmopsheric ice nucleation by mineral dust  (?, ?). Not surprisingly, the exceptional ice nucleation efficiency of K-feldspar has garnered substantial research interest over the last decade, driving extensive investigations aimed at elucidating the mechanisms of ice nucleation and identifying the nature of the active sites responsible for it  (?, ?, ?, ?, ?, ?, ?, ?). On the micron scale, pioneering work by Kiselev et al. (?) and subsequent studies by others (?, ?) conclusively showed that defects such as steps, cracks, and cavities are responsible for the high ice nucleation efficiency of feldspar particles, based on direct observations using optical microscopy. Furthermore, a systematic and reproducible orientation of the ice crystals with respect to the feldspar substrate was observed, highlighting that a specific crystallographic plane of feldspar must govern ice nucleation  (?). In Fig. 1A, we show schematically how defects, such as steps, expose patches of different crystallographic surfaces, which can serve as active sites for ice nucleation. However, at the atomic scale, the mechanism remains controversial and current hypothesis are the subject of significant debate. Molecular simulations based on energy minimization that accompany the experimental work of Kiselev et al.  (?) suggest that the primary prismatic plane of hexagonal ice forms on the K-feldspar (100) surface. In stark contrast, extensive molecular dynamics (MD) simulations by Soni and Patey did not find evidence of high ice nucleation efficiency at the (100), (010), and (001) surfaces of feldspar  (?). Moreover, recent experimental work based on atomic force microscopy has been restricted to the easily cleavable (001) surface and no indication of high ice nucleation efficiency has been found  (?, ?). In spite of the progress described above, the most likely mechanism by which ice forms in our planet’s atmosphere remains unknown.
Figure 1: K-feldspar serves as an effective ice-nucleating particle in the atmosphere owing to structural defects that promote ice nucleation. (A) Schematic illustration of the heterogeneous ice nucleation at K-feldspar defects with exposed non-perfect cleavage planes. The pink and green dashed lines highlight the first and second prismatic surfaces of ice, which are approximately parallel to the (100) and (110) planes of feldspar, respectively. (B) The triclinic crystal structure of microcline K-feldspar. (C) Crystallographic planes of K-feldspar: (001), (010), (100), (110), ($\bar{1}$10), and ($\bar{2}$01). (D) Configuration for molecular dynamics simulations with one termination of the K-feldspar (110) surface in contact with water above.
In this work, we use MD simulations to study a broad range of K-feldspar surfaces exposed to water, including the (001), (010), (100), (110), ($\bar{1}$10), and ($\bar{2}$01) surfaces. The interatomic forces driving these simulations are derived from a machine-learning interatomic potential (MLP) trained to achieve a very high accuracy using quantum-mechanical electronic-structure theory calculations. Our simulations reveal that the K-feldspar (110) surface structures interfacial liquid water into an arrangement closely resembling that on the (110) surface of cubic ice (equivalent to the secondary prismatic surface of hexagonal ice). We also carried out enhanced sampling molecular simulations in order to bring ice nucleation, which is a rare event in the timescale of standard MD simulations, within reach. The ice clusters that nucleate at the (110) surface during these simulations exhibit a cubic-ice structure, with the (110) plane aligned with the (110) K-feldspar surface. In contrast, none of the other feldspar surfaces show signatures of high ice nucleation potency. Our results shed new light on the molecular mechanism for heterogeneous ice nucleation on K-feldspar, which is the main route for ice formation in the atmosphere.
Structure of water at multiple surfaces
K-feldspar (KAlSi3O8) is a tectosilicate mineral characterized by a three-dimensional framework of corner-sharing SiO4 and AlO4 tetrahedra, with K+ occupying interstitial sites. We studied the most stable polymorph of K-feldspar (microcline) at room temperature, which has a triclinic crystal structure (Fig. 1B). Microcline shows perfect cleavage along the (001) plane and good cleavage along the (010) plane, and less frequently is also known to expose the (100), (110), ($\bar{1}$10), and ($\bar{2}$01) planes. These planes are shown in Fig. 1C. We systematically explored candidate surfaces and identified a total of 13 distinct surface terminations. Further details about the methodology and a description of the surface terminations is provided in the Methods section and in Tables S1 and S2 of the Supplementary Material (SM). We then constructed large interfacial configurations for all terminations of K-feldspar in contact with water to investigate the behavior of the candidate surfaces using MD simulations. A representative configuration of the (110)-$\alpha$ surface is shown in Fig. 1D. In order to describe the interactions in this system with first-principles accuracy, we trained an MLP using highly accurate quantum-mechanical electronic structure calculations for all surface terminations discussed above (see Methods for details), and we used this MLP to drive MD simulations. One of us recently leveraged this technique to study homogeneous ice nucleation and found that nucleation rates are predicted in good agreement with experiment  (?).
[…]
To identify the feldspar surfaces that promote ice nucleation among the 13 candidate terminations, we investigated the structure of interfacial water, which is known to be a useful descriptor of the ice nucleation potency of a surface  (?). First, we analyzed the water density along the direction perpendicular to the feldspar surfaces. In Fig. 2A, we show the water density profile at one particular surface, namely, the feldspar (110)-$\alpha$ surface. It is reasonable to expect that a water density similar to that on ice would facilitate ice nucleation. For this reason, we compared our results with the water density on the surfaces of ice polymorphs that can form under Earth’s atmospheric conditions, namely hexagonal ice (Ih) and cubic ice (Ic). As an example, we show in Fig. 2B the water density profile at the ice Ic (110) surface. To facilitate a clear comparison, we calculated the difference between the two density profiles, as shown in Fig. 2C. A small difference indicates that the water density at a given $z$ position is similar for the feldspar and ice surfaces. Water density profiles for other feldspar and ice surfaces, as well as their difference, are presented in Fig. S1 and S2. Furthermore, to quantitatively evaluate the overall deviation of water densities at the feldspar–water interfaces with respect to the densities at ice–water interfaces, we calculated the mean squared error, $\mathrm{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left[\rho_{\mathrm{felds}}(z_{i})-\rho_{\mathrm{ice}}(z_{i})\right]^{2},$ where $\rho_{\mathrm{felds}}(z_{i})$ and $\rho_{\mathrm{ice}}(z_{i})$ denote the water densities at the feldspar–water and ice-water interfaces, respectively, at position $z_{i}$, and $N$ is the total number of sampling points within $z$ $\leq$} 15 Å. In Fig. 2D, we show the MSE between water densities at the 13 feldspar surfaces, and at ice surfaces of ice Ih and ice Ic. Feldspar surfaces show a variety of behaviors, with some of them being very different from all ice surfaces, while a few of them give rise to water density profiles that closely match the water density on one of the studied ice surfaces. We find that the minimum MSE corresponds to the feldspar (110)-$\alpha$ surface and the ice Ic (110) surface, indicating that the interfacial water densities at these two surfaces are nearly identical (Fig. 2A and B). There are also a few other feldspar surfaces with low MSE, yet the feldspar (110)-$\alpha$ surface stands out as the most promising candidate. Note that the feldspar (100)-$\alpha$ surface, previously proposed as the active site for ice nucleation (?), shows a very high MSE with respect to all ice surfaces, and thus the structure of water at this surface does not show signatures of high ice nucleation potency.
We now turn to a detailed analysis of the structure of water at the feldspar (110)-$\alpha$ surface. The density profile of water at this surface exhibits several peaks near the interface (see Fig. 2A). The first peak is notably high, indicating that water molecules in this region are highly structured, and the height, width, and spacing between other density peaks closely match those of the water density profile on the ice Ic (110) surface (see Fig. 2B). Furthermore, we examined the water-density isosurface (70 molecules/nm3) in the xxyy plane corresponding to the first peak. As shown in Fig. 2E, water molecules within this first peak tend to reside near hydroxyl groups on the feldspar surface, forming a pattern with an irregular hexagon shape that closely resembles that observed on the ice Ic (110) surface (Fig. 2F). This atomic arrangement strongly suggests that ice could readily grow on this surface, with the nascent ice Ic crystal oriented such that its (110) plane is aligned with the feldspar (110) surface. Note that the distribution of hydroxyl groups at the feldspar (110)-$\alpha$ surface, shown in Fig. 2G, does not fully match the distribution of water molecules at the ice Ic (110) surface. In spite of the absence of a perfect lattice match, the feldspar (110)-$\alpha$ surface is still able to create a distribution of interfacial water which closely matches the distribution of water at the ice Ic (110) surface. This phenomenon has already been observed in other minerals, such as mica and kaolinite  (?, ?), and represents a different point of view with respect to the early work on heterogeneous ice nucleation by Vonnegut  (?), which emphasized the role of lattice match. We also note that the feldspar (110) surface is one of the three candidate planes that can explain the orientation of ice crystals relative to feldspar observed by Kiselev et al.  (?) using scanning electron microscopy. We analyzed the two other possible feldspar planes compatible with the data of Kiselev et al., namely, the (100) and the (010) planes, and we find, in Fig. S1 and S3, that the density profiles of interfacial water and its in-plane distribution does not match the arrangements found on any ice surface.
Taken together, the results described above strongly suggest that the (110)-$\alpha$ surface of feldspar is responsible for its high ice nucleation potency. Below, we shall investigate this hypothesis in further detail.
Formation of ice at the feldspar (110) surface
Figure 3: Formation of ice at the feldspar (110)-$\alpha$ surface. (A) Snapshots of ice nucleation on the feldspar surface over the $Q_{6}$ order parameter from MD simulations with a bias along the $Q_{6}$. (B) Number of ice-like water molecules and $Q_{6}$ versus simulation time. (C) Water-density isosurfaces of the first ice layer at 70 molecules/nm3. (D) Ice-like water molecules over time at different supercooling, ${\Delta}T$. (E) Critical cluster size, $N^{*}$, versus ${\Delta}T$ with a classical nucleation theory (CNT) fit (dashed lines). Data for homogeneous nucleation is taken from Ref.  (?). (F) Interfacial free energy, $\gamma$, versus ${\Delta}T$. The dashed lines represent a fit based on CNT. (G) Solid lines show the CNT-calculated free-energy barriers for heterogeneous (${\Delta}G^{*}_{\mathrm{het}}$) on the feldspar surface and homogeneous (${\Delta}G^{*}_{\mathrm{hom}}$) nucleation. The heterogeneous nucleation data point (${\Delta}G^{*}_{\mathrm{het-US}}$) comes from umbrella sampling (panel H), and the homogeneous point is from Ref.  (?). (H) Free-energy profile along $Q_{6}$ obtained from umbrella sampling simulations. (I) Heterogeneous nucleation rates ($J$) versus ${\Delta}T$ with CNT fit and experimental estimates. Further details of the ice-structure identification, CNT calculations, and estimation of nucleation rates are provided in the Methods section.
We now seek to observe directly the ice nucleation process on the (110)-$\alpha$ feldspar surface and elucidate its mechanism. This cannot be achieved using standard MD simulations due to the high free energy barrier for this process, and thus we resort to the application of enhanced sampling techniques, which bring the nucleation process within affordable simulation time. In particular, we drive the formation of ice using steered MD simulations with the $Q_{6}$ Steinhardt order parameter, which is an effective and widely-used collective variable (order parameter) to study crystal nucleation  (?). This methodology is based on introducing a harmonic bias potential, as a function of $Q_{6}$ order parameter, whose center moves from a $Q_{6}$ value corresponding to liquid water to a $Q_{6}$ value corresponding to ice, thereby promoting crystallization (see details in the Methods section).
In Fig. 3A, we show that during the simulation ice nucleates from the interface and eventually forms an approximately hemispherical ice cluster on the surface, as the $Q_{6}$ value increases under the applied bias potential (Fig. 3B). Figure 3B shows that the cluster has a structure compatible with ice Ic, instead of ice Ih (see Methods for further information about this analysis). We also investigated the arrangement of water molecules in the interfacial layer and found that it remains largely unchanged before and after ice nucleation, although the distribution becomes more structured after ice is formed (compare Fig. 3C and Fig. 2E). This result is highly robust, with ice Ic consistently formed in multiple independent simulations regardless of the applied biasing rate (see Fig. S4). In contrast, simulations on the (010), (100) and other surfaces do not show ice nucleation and growth from the interface and, instead, ice nucleates from the bulk water region, indicating that these surfaces cannot effectively promote ice nucleation (see Fig. S5 and S6).
Next, we investigate the thermodynamics of ice nucleation, including the calculation of the critical cluster size and free energy barrier for this process. Starting from the ice cluster formed on the feldspar (110)-$\alpha$ surface and immersed in water, we performed seeding simulations. This methodology is based on carrying out standard MD simulations starting from ice seeds of various sizes at different temperatures (see Fig. 3D and Fig. S7). For each seed size, we study whether the ice cluster grows or melts at a given temperature and use this data to determine the temperature at which a seed is critical, i.e., it has the same probability of growing and melting. We show in Fig. 3E the relationship between the temperature and the critical cluster size ($N^{*}_{\mathrm{het}}$), measured as the number of ice-like water molecules. In Fig. 3E, we also show the critical cluster sizes in homogeneous nucleation ($N^{*}_{\mathrm{hom}}$) (?), and it can be observed that for a given supercooling, the critical cluster size at the (110)-$\alpha$ feldspar surface is between 2 and 4 times smaller than in the homogeneous nucleation case. Note that the melting temperature in our SCAN-based MLP is $T_{m}=308\ \rm K$ (see Fig. S8), which agrees with previous simulations but is higher than the experimental value  (?, ?, ?), and therefore we express all results in terms of supercooling ${\Delta}T=T_{m}-T$.
With the critical cluster sizes computed above, we can obtain thermodynamic properties using the organizing framework of classical nucleation theory (see Methods section). In Fig. 3F, we show the effective interfacial free energy $\gamma_{\mathrm{eff}}$ for the heterogeneous nucleation process and the water-ice interfacial free energy $\gamma_{\mathrm{liq-ice}}$, relevant for homogeneous nucleation, which was reported in ref. ?. These two quantities are related by $\gamma_{\mathrm{eff}}=\gamma_{\mathrm{liq-ice}}f(\theta)$, where $f(\theta)<1$ is a function of the contact angle $\theta$  (?). In the studied temperature range, $\gamma_{\mathrm{eff}}$ is lower than $\gamma_{\mathrm{liq-ice}}$ by around 30%, indicating that the (110)-$\alpha$ surface indeed is a suitable substrate for heterogeneous ice nucleation. Using classical nucleation theory, we also computed the nucleation free energy barrier and the results are shown in Fig. 3G. The heterogeneous nucleation free-energy barrier (${\Delta}G^{*}_{\mathrm{het}}$) is significantly lower than that of homogeneous ice nucleation (${\Delta}G^{*}_{\mathrm{hom}}$), showing the increased likelihood of observing a fluctuation at the surface that reaches the critical cluster size. We also computed the free energy barrier using umbrella sampling calculations (Fig. 3H) and the results are in good agreement with the estimates from classical nucleation theory. Using the results described above, we computed the heterogeneous nucleation rate ($J$) and, as we shall see, it provides a direct link with experiment. In Fig. 3I, we show $J$ as a function of temperature and we compare it with an estimate of the nucleation rates based on the number of nuclei per unit area in optical microscopy experiments  (?). For such nucleation rates, our curve predicts a heterogeneous nucleation temperature for the (110)-$\alpha$ feldspar of 23 K, in very good agreement with the estimates of Atkinson et al. based on the droplet-freezing technique  (?).
Orientation relationship between ice and feldspar
Figure 4: Orientation relationship between feldspar and the ice cluster formed in our simulation. (A) Orientation and lattice of ice formed on the K-feldspar (110)-$\alpha$ surface. Visualizations of the ice structure along the $x$, $y$, and $z$ axis are shown. Characteristic patterns on each ice plane are depicted using black dashed lines. (B) Illustration of the ice basal-plane axis orientation relative to the feldspar (001), (010), and (110) planes. Note the basal (0001) plane of hexagonal ice (Ih) is equivalent to the (111) plane of cubic ice (Ic).
[…]
Conclusions
Our MD simulations show that the (110) surface of K-feldspar is responsible for its high ice nucleation efficiency. A comprehensive analysis of alternative feldspar surfaces shows no evidence of enhanced ice nucleation. Our results also provide two key molecular-level insights into this phenomenon. First, we find that the high ice nucleation efficiency of the (110) feldspar surface is explained by its unique ability to structure liquid water into an arrangement closely resembling that on the (110) surface of cubic ice, thereby providing an optimal template for ice nucleation and growth. Second, the ice clusters formed over the (110) feldspar surface are approximately hemispherical, with minor faceting, and mostly composed of cubic ice, a metastable polymorph with respect to hexagonal ice. These findings stand in contrast to previous work, which had suggested that the high-energy feldspar (100) surface provides preferential sites for ice nucleation. Our results contradict this hypothesis and instead support a new interpretation in which atmospheric heterogeneous ice nucleation occurs at feldspar defects exposing patches of the (110) surface. In this new picture, the initial nuclei adopt a cubic ice structure but subsequently transform into hexagonal ice as they grow to macroscopic dimensions. Our work provides a clear molecular-level explanation for the exceptionally high ice nucleation efficiency of K-feldspar, which has been shown to be the dominant ice nucleating particle in the atmosphere.
Acknowledgments
We are grateful for the computational resources provided on the Hyperion cluster at the Donostia International Physics Center (DIPC) and on MareNostrum 5 ACC at the Barcelona Supercomputing Center, granted through the Spanish Supercomputing Network (RES) allocations FI-2024-2-0026 and FI-2024-3-0028. We thank Prof. Hu Qiu from Nanjing University of Aeronautics and Astronautics for valuable suggestions on the data presentation in this work.
Funding:
W.Z. was funded by the ’María de Maeztu’ Units of Excellence program No. CEX2020-001038-M. P.M.P. acknowledges funding from the Marie Sklodowska-Curie Cofund Programme of European Commission project H2020-MSCA-COFUND-2020-101034228-WOLFRAM2.
Author contributions:
W.Z. performed research. W.Z. and P.M.P. designed research, analyzed data, and wrote the manuscript. P.M.P conceived the project.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
The scripts used for generating slab surfaces, the dataset for training the MLPs, the input files for MLP training, the final trained MLP, as well as the input files for the MD simulations and DFT calculations reported in this work, are openly available on Zenodo (DOI https://doi.org/10.5281/zenodo.17610332)  (?).
Supplementary materials
Methods
Tables S1 to S2
Figs. S1 to S13
References (30-53)