### v3

# Molecular mechanism of heterogeneous ice nucleation on potassium feldspar Wanqi Zhou1, and

Abstract
Mineral dust aerosols strongly influence Earth’s climate by acting as ice-nucleating particles (INPs). Feldspar minerals, particularly K-feldspar, are recognized as key INPs, and a previous study attributed this behavior to (100) surfaces exposed at defects. Using machine-learning molecular dynamics simulations, we systematically investigate ice nucleation on multiple K-feldspar surfaces. We identify the (110) surface, exposed at defects such as steps, as the most active plane for ice formation. This surface uniquely structures interfacial water into an arrangement resembling that on the (110) surface of cubic ice, providing an optimal template for nucleation. Using advanced sampling, we directly observe the formation of clusters with cubic-ice structure and their orientation agrees with experiment. These results provide molecular-level insight into one of the most important ice nucleation processes in our planet’s atmosphere.
Introduction
Ice nucleation in the atmosphere has an important impact on the climate of our planet, affecting precipitation formation, the hydrological cycle, and Earth’s radiation budget[29, 52]. Understanding the mechanism of this process is essential for climate modeling and prediction, a subject of increasing relevance due to anthropogenic climate change[7], as well as for multiple applications such as cloud seeding, the production of artificial snow, anti-icing strategies for aircraft, and frost protection in agriculture. Ice can nucleate from liquid water via two distinct pathways[24]. The first is homogeneous ice nucleation, which occurs in pure supercooled water at sufficiently low temperatures, namely, below about -38°C. More commonly, ice forms via heterogeneous ice nucleation, where it nucleates on a wide variety of atmospheric aerosols known as ice-nucleating particles (INPs)[24, 17]. These particles provide sites or surfaces where ice can readily form, thereby significantly reducing the free energy barrier for the transformation[19]. As a result, heterogeneous ice nucleation can occur at warmer conditions, typically between -38°C and -10°C[17].
Several types of INPs are known to contribute to atmospheric ice nucleation[20, 36]. Mineral dust particles are considered the globally dominant INPs, owing to their relatively high ice-nucleating ability and abundance[20, 3]. However, other types of INPs can also play an important role, and recent evidence highlights the relevance of organic aerosols, particularly in regions far from major dust sources[57, 53, 15]. Until recently, the specific components responsible for the ice-nucleating activity of mineral dusts was not fully understood. A major breakthrough came in 2013[1, 58], when it was shown that potassium (K)-feldspar dominates ice nucleation by mineral dust[14, 13, 15], and thus represents one of the primary pathways for ice formation in the atmosphere. Given its exceptional ice nucleation efficiency, K-feldspar has garnered substantial research interest over the last decade, driving extensive investigations aimed at elucidating the mechanisms of ice nucleation and identifying the nature of the active sites responsible for it[23, 56, 16, 39, 31, 22, 25, 21]. On the micron scale, pioneering work by Kiselev et al.[23] and subsequent studies by others[31, 16] conclusively showed that defects such as steps, cracks, and cavities are responsible for the high ice nucleation efficiency of feldspar particles, based on direct observations using scanning electron microscopy. Furthermore, a systematic and reproducible orientation of the ice crystals with respect to the feldspar substrate was observed, highlighting that a specific crystallographic plane of feldspar must govern ice nucleation[23]. In Fig. 1A, we show schematically how defects, such as steps, expose patches of different crystallographic surfaces, which can serve as active sites for ice nucleation. However, at the atomic scale, the mechanism remains controversial and current hypotheses are the subject of significant debate. Molecular simulations based on energy minimization that accompany the experimental work of Kiselev et al.[23] suggest that the primary prismatic plane of hexagonal ice forms on the K-feldspar (100) surface. In stark contrast, extensive molecular dynamics (MD) simulations by Soni and Patey did not find evidence of high ice nucleation efficiency at the (100), (010), and (001) surfaces of feldspar[39]. Moreover, recent experimental work based on atomic force microscopy has been restricted to the easily cleavable (001) surface and no indication of high ice nucleation efficiency has been found[8, 5]. In spite of the progress described above, the mechanism underlying one of the most important ice nucleation processes in our planet’s atmosphere remains unknown.
In this work, we use MD simulations to study a broad range of K-feldspar surfaces exposed to water, including the (001), (010), (110), (100), (10), and (01) surfaces. The interatomic forces driving these simulations are derived from a machine-learning interatomic potential (MLP) trained to achieve a very high accuracy using quantum-mechanical electronic-structure theory calculations. Our simulations reveal that the K-feldspar (110) surface structures interfacial liquid water into an arrangement closely resembling that on the (110) surface of cubic ice (equivalent to the secondary prismatic surface of hexagonal ice). We also carried out enhanced sampling molecular simulations in order to bring ice nucleation, which is a rare event in the timescale of standard MD simulations, within reach. The ice clusters that nucleate at the (110) surface during these simulations exhibit a cubic-ice structure, with the (110) plane aligned with the (110) K-feldspar surface. In contrast, none of the other feldspar surfaces show signatures of high ice nucleation potency. Our results shed new light on the molecular mechanism for heterogeneous ice nucleation on K-feldspar, which is a key route for ice formation in the atmosphere.
Results and Discussion
Structure of water at multiple surfaces
K-feldspar (KAlSi3O8) is a tectosilicate mineral characterized by a three-dimensional framework of corner-sharing SiO4 and AlO4 tetrahedra, with K+ occupying interstitial sites. We studied microcline, which is the most stable polymorph of K-feldspar at room temperature and has a triclinic crystal structure (Fig. 1B). Microcline shows perfect cleavage along the (001) plane and good cleavage along the (010) plane, and less frequently is also known to expose the (110), (100), (10), and (01) planes. These planes are shown in Fig. 1C. We systematically explored candidate surfaces and identified a total of 13 distinct surface terminations. Further details about the methodology and a description of the surface terminations is provided in the Methods section and in Tables S1 and S2 of the Supplementary Material (SM). We then constructed large interfacial configurations for all terminations of K-feldspar in contact with water to investigate the behavior of the candidate surfaces using MD simulations. A representative configuration of the (110)- surface is shown in Fig. 1D. In order to describe the interactions in this system with first-principles accuracy, we trained an MLP using quantum-mechanical electronic structure calculations for all surface terminations discussed above (see Methods for details), and we used this MLP to drive MD simulations. One of us recently leveraged this technique to study homogeneous ice nucleation and found that nucleation rates are predicted in good agreement with experiment[34].
To identify the feldspar surfaces that promote ice nucleation among the 13 candidate terminations, we investigated the structure of interfacial water, which is known to be a useful descriptor of the ice nucleation potency of a surface[42]. First, we analyzed the water density along the direction perpendicular to the feldspar surfaces. In Fig. 2A, we show the water density profile at one particular surface, namely, the feldspar (110)- surface. It is reasonable to expect that a water density similar to that on ice would facilitate ice nucleation. For this reason, we compared our results with the water density on the surfaces of ice polymorphs that can form under Earth’s atmospheric conditions, namely hexagonal ice (Ih) and cubic ice (Ic). As an example, we show in Fig. 2B the water density profile at the ice Ic (110) surface. To facilitate a clear comparison, we calculated the difference between the two density profiles, as shown in Fig. 2C. A small difference indicates that the water density at a given position is similar for the feldspar and ice surfaces. Water density profiles for other feldspar and ice surfaces, as well as their difference, are presented in Fig. S1 and S2. Furthermore, to quantitatively evaluate the overall deviation of water densities at the feldspar–water interfaces with respect to the densities at ice–water interfaces, we calculated the mean squared error, where and denote the water densities at the feldspar–water and ice-water interfaces, respectively, at position , and is the total number of sampling points within 15 Å. In Fig. 2D, we show the MSE between water densities at the 13 feldspar surfaces, and at ice surfaces of ice Ih and ice Ic. Feldspar surfaces show a variety of behaviors, with some of them being very different from all ice surfaces, while a few of them give rise to water density profiles that closely match the water density on one of the studied ice surfaces. We find that the minimum MSE corresponds to the feldspar (110)- surface and the ice Ic (110) surface, indicating that the interfacial water densities at these two surfaces are nearly identical (Fig. 2A and B). There are also a few other feldspar surfaces with low MSE, yet the feldspar (110)- surface stands out as the most promising candidate. Note that the feldspar (100)- surface, previously proposed as the active site for ice nucleation[23], shows a very high MSE with respect to all ice surfaces, and thus the structure of water at this surface does not show signatures of high ice nucleation potency.
We now turn to a detailed analysis of the structure of water at the feldspar (110)- surface. The density profile of water at this surface exhibits several peaks near the interface (see Fig. 2A). The first peak is notably high, indicating that water molecules in this region are highly structured, and the height, width, and spacing between other density peaks closely match those of the water density profile on the ice Ic (110) surface (see Fig. 2B). Furthermore, we examined the water-density isosurface (70 molecules/nm3) in the – plane corresponding to the first peak. As shown in Fig. 2E, water molecules within this first peak tend to reside near hydroxyl groups on the feldspar surface, forming a pattern with an irregular hexagon shape that closely resembles that observed on the ice Ic (110) surface (Fig. 2F). This pattern does not match the arrangement of water molecules on any other ice surface. In particular, the pattern does not resemble the water distribution at the (110) plane of ice Ih, equivalent to the (110) plane of ice Ic, suggesting a strong polymorph selectivity (see Fig. S3). The arrangement of water molecules at feldspar (110)- strongly suggests that ice could readily grow on this surface, with the nascent ice Ic crystal oriented such that its (110) plane is aligned with the feldspar (110) surface. Note that the distribution of hydroxyl groups at the feldspar (110)- surface, shown in Fig. 2G, does not fully match the distribution of water molecules at the ice Ic (110) surface. In spite of the absence of a perfect lattice match, the feldspar (110)- surface is still able to create a distribution of interfacial water which closely matches the distribution of water at the ice Ic (110) surface. This phenomenon has already been observed in other minerals, such as mica and kaolinite[41, 40], and represents a different point of view with respect to the early work on heterogeneous ice nucleation by Vonnegut[54], which emphasized the role of lattice match. We also note that the feldspar (110) surface is one of the three candidate planes that can explain the orientation of ice crystals relative to feldspar observed by Kiselev et al.[23] using scanning electron microscopy. We analyzed the two other possible feldspar planes compatible with the data of Kiselev et al., namely, the (100) and the (010) planes, and we find, in Figs. 2D, S1 and S4, that the density profiles of interfacial water and its in-plane distribution does not match the arrangements found on any ice surface.
Taken together, the results described above strongly suggest that the (110)- surface of feldspar is responsible for its high ice nucleation potency. Below, we shall investigate this hypothesis in further detail.
Formation of ice at the feldspar (110) surface
We now seek to observe directly the ice nucleation process on the (110)- feldspar surface and elucidate its mechanism. This cannot be achieved using standard MD simulations due to the high free energy barrier for this process, and thus we resort to the application of enhanced sampling techniques, which bring the nucleation process within affordable simulation time. In particular, we drive the formation of ice using steered MD simulations with the Steinhardt order parameter, which is an effective and widely-used collective variable (order parameter) to study crystal nucleation[43]. This methodology is based on introducing a harmonic bias potential, as a function of the order parameter, whose center moves from a value corresponding to liquid water to a value corresponding to ice, thereby promoting crystallization (see details in the Methods section).
In Fig. 3A, we show that during the simulation ice nucleates from the interface and eventually forms an approximately hemispherical ice cluster on the surface, as the value increases under the applied bias potential. Figure 3B shows that the cluster has a structure compatible with ice Ic, instead of ice Ih. This result is highly robust, with ice Ic consistently formed in multiple independent simulations regardless of the applied biasing rate (see Fig. S5). Although the introduction of a bias potential could affect polymorph selection, the results of our biased simulations lend additional support to our previous finding that the (110)- feldspar surface templates the formation of ice Ic. We also investigated the arrangement of water molecules in the interfacial layer and found that it remains largely unchanged before and after ice nucleation. Once ice is formed, the distribution of water becomes more structured and almost identical to that on the (110) ice Ic surface (compare Fig. 3C and Fig. 2E). Thus, both our standard (unbiased) and biased simulations of water at the (110)- feldspar surface support a strong preference for the formation of ice Ic. In contrast, simulations of the (100) and other feldspar surfaces do not show ice nucleation and growth from the interface and, instead, ice nucleates from the bulk water region, indicating that these surfaces cannot effectively promote ice nucleation (see Fig. S6 and S7).
Next, we investigate the thermodynamics of ice nucleation, including the calculation of the critical cluster size and free energy barrier for this process. Starting from the ice cluster formed on the feldspar (110)- surface and immersed in water, we performed seeding simulations. This methodology is based on carrying out standard MD simulations starting from ice seeds of various sizes at different temperatures (see Fig. 3D and Fig. S8). For each seed size, we study whether the ice cluster grows or melts at a given temperature and use this data to determine the temperature at which a seed is critical, i.e., it has the same probability of growing and melting. We show in Fig. 3E the relationship between the temperature and the critical cluster size (), measured as the number of ice-like water molecules. In Fig. 3E, we also show the critical cluster sizes in homogeneous nucleation ()[34], and it can be observed that for a given supercooling, the critical cluster size at the (110)- feldspar surface is between 2 and 4 times smaller than in the homogeneous nucleation case. Note that the melting temperature in our MLP is (see Fig. S9), which agrees with previous simulations but is higher than the experimental value[34, 33, 4], and therefore we express all results in terms of supercooling .
With the critical cluster sizes computed above, we can obtain thermodynamic properties using the organizing framework of classical nucleation theory (see Methods section). In Fig. 3F, we show the effective interfacial free energy () for the heterogeneous nucleation process and the water-ice interfacial free energy (), relevant for homogeneous nucleation, which was reported in ref. Piaggi et al. [34]. These two quantities are related by , where is a function of the contact angle [19]. In the studied temperature range, is lower than by around 30%, indicating that the (110)- surface indeed is a suitable substrate for heterogeneous ice nucleation. Using classical nucleation theory, we also computed the nucleation free energy barrier and the results are shown in Fig. 3G. The heterogeneous nucleation free-energy barrier () is significantly lower than that of homogeneous ice nucleation (), showing the increased likelihood of observing a fluctuation at the surface that reaches the critical cluster size. We also computed the free energy barrier using umbrella sampling calculations (Fig. 3H) and the results are in good agreement with the estimates from classical nucleation theory. Using the results described above, we computed the heterogeneous nucleation rate () and, as we shall see, it provides a direct link with experiment. In Fig. 3I, we show as a function of temperature and we compare it with an estimate of the nucleation rates based on the number of nuclei per unit area in microscopy experiments[23]. For such nucleation rates, our curve predicts a heterogeneous nucleation temperature for the (110)- feldspar surface of 23 K, in very good agreement with the estimates of Atkinson et al. based on the droplet-freezing technique[1]. This temperature is characteristic of the active sites that lead to the typical ice nucleation behavior of feldspar, which are the sites studied by Kiselev et al.[23]. We note that other experiments report much higher nucleation temperatures of around -2°C for extremely rare yet highly active sites, which due to their low abundance, have a limited relevance for ice nucleation at atmospheric conditions[14].
Orientation relationship between ice and feldspar
From the simulations reported above, we can analyze the orientation of ice formed on the feldspar (110)- surface and compare the results with experimental observations. By direct visual inspection, we confirmed that the (110) plane of ice Ic is parallel to the feldspar (110)- surface (see Fig. 4A), as we had already inferred from the results in Fig. 2. This result is fully compatible with the experimental data of Kiselev et al.[23], which can be explained by the (110) plane of ice Ih, equivalent to the ice Ic (110) plane (see the structural equivalence in Fig. S10), aligned with the (110) feldspar surface. Furthermore, we show in Fig. 4A that the ice Ic (111) plane is aligned with the plane (perpendicular to the axis). Using this information, we computed an angle of between the ice Ic (111) plane and the feldspar (001) plane, as shown in Fig. 4B. This result agrees very well with the measurements of Kiselev et al. that found an angle of approximately between the basal axis of ice Ih (corresponding to the (111) plane of ice Ic) and the feldspar (001) surface[23]. Therefore, the orientation of ice found in our simulation is fully compatible with all the available experimental evidence. A more thorough analysis of the orientation of ice Ic with respect to feldspar is presented in the Fig. S11. We note that experiments consistently show the formation of ice crystals with hexagonal shape, which are the hallmark of ice Ih, while our simulations predict the formation of a cluster of ice Ic at this surface. This apparent discrepancy is resolved by noting that microscopic ice clusters may consist of cubic or stacking-faulted ice, but as they grow towards macroscopic sizes they must transform to ice Ih, which is the equilibrium polymorphic form.
Conclusion
Our MD simulations show that the (110) surface of K-feldspar is responsible for its high ice nucleation efficiency. A comprehensive analysis of alternative feldspar surfaces shows no evidence of enhanced ice nucleation. Our results also provide two key molecular-level insights into this phenomenon. First, we find that the high ice nucleation efficiency of the (110) feldspar surface is explained by its unique ability to structure liquid water into an arrangement closely resembling that on the (110) surface of cubic ice, thereby providing an optimal template for ice nucleation and growth. Second, the ice clusters formed over the (110) feldspar surface are approximately hemispherical, with minor faceting, and mostly composed of cubic ice, a metastable polymorph with respect to hexagonal ice. These findings stand in contrast to previous work, which had suggested that the high-energy feldspar (100) surface provides preferential sites for ice nucleation. Our results contradict this hypothesis and instead support a new interpretation in which atmospheric heterogeneous ice nucleation occurs at feldspar defects exposing patches of the (110) surface. In this new picture, the initial nuclei adopt a cubic ice structure but subsequently transform into hexagonal ice as they grow to macroscopic dimensions. Although the proposed mechanism relies on the exposure of the (110) plane at defect sites, we note that this surface is common and relatively easy to cleave, ranking just below the (001) and (010) surfaces in terms of cleavage perfection[38]. Our work provides a clear molecular-level explanation for the exceptionally high ice nucleation efficiency of K-feldspar, which has been shown to be one of the key ice nucleating particles in the atmosphere.
Methods
Generating feldspar surface terminations
All possible terminations of the (001), (010), (100), (110), (10), and (01) planes were generated using the SlabGenerator module of the pymatgen package[49]. A cutoff of 1 Å along the surface normal direction was applied to determine whether atoms lie within the same atomic plane, which was used to filter the generated terminations. There is extensive experimental evidence for the absence of reconstructions in feldspar surfaces and thus we only consider flat unreconstructed surfaces[8, 5, 6]. Hydrogen atoms were inserted near undercoordinated oxygen atoms at the surfaces, in accordance with recent atomic force microscopy observations, which show full surface hydroxylation[8]. In addition, surface terminations that did not preserve the integrity of SiO4 and AlO4 polyhedra or failed to maintain charge neutrality were discarded. Consequently, a total of 13 distinct surface terminations were obtained, including 2 terminations for the (001) plane, 2 for the (010) plane, 2 for the (100) plane, 3 for the (110) plane, 3 for the (10) plane, and 1 for the (01) plane (see Table S1 and S2). Among these, all terminations of the (001), (010), and (100) planes correspond to those previously reported, namely, the (001)-, (001)-, (010)-, (010)-, (100)-, and (100)- surfaces[39, 33]. Three terminations were identified for the (110) plane, each of them involving the breaking of eight bonds per unit cell (Table S1). In one of the terminations, three oxygen atoms of a SiO4 tetrahedron are disconnected from their neighboring two Si and one Al atoms, leaving only one oxygen atom bonded to the surface. This configuration is likely unstable and was thus denoted as the (110)- termination. In another case, the cleavage requires breaking three bonds of a SiO4 tetrahedron, which is energetically more unfavorable than cases where only one or two bonds are broken. This termination was therefore labeled as (110)-, while the remaining, more stable termination was designated as (110)-. The (10) plane is a mirror-related counterpart of the (110) plane, and consequently exhibits atomic-scale structural features similar to those of the (110) surface. Therefore, the terminations of the (10) plane were named following the same convention, namely, (10)-, (10)-, and (10)-. The (01) plane has only a single termination, which is therefore designated as (01)-.
The large K-feldspar slabs used for molecular dynamics (MD) simulations expose two identical surfaces along the -axis, with atomic-scale structures of both surfaces corresponding to one of the terminations listed above. One of the surfaces is in direct contact with water, forming the water–feldspar interface. A vacuum layer is introduced along the -axis between the water layer and the opposite surface of the feldspar slab to eliminate spurious interactions.
Electronic structure calculations
Plane-wave DFT calculations were performed using the Quantum ESPRESSO suite for electronic structure calculations v6.4.1[10, 9]. We used the Strongly Constrained and Appropriately Normed (SCAN) exchange and correlation functional[46, 45] as implemented in the LIBXC 4.3.4 library[28]. We employed norm-conserving, scalar-relativistic pseudopotentials[12] for K, Al, Si, O, H parametrized using the PBE[32] functional with 9, 11, 4, 6, and 1 valence electrons, respectively. Kinetic energy cutoffs of 110 and 440 Ry were used for the wave functions and the charge density. We used only the -point in our calculations. All other parameters were set to their default values in Quantum ESPRESSO[10, 9].
Machine learning interatomic potential
The smooth-edition of the Deep Potential methodology developed by Zhang et al., as implemented in DeePMD-kit v2.10.0, was used to train the machine-learning interatomic potentials (MLPs)[59, 55, 60]. An active learning strategy was employed during the training process, as shown in Fig. S12. First, MD simulations driven by our previous MLP (Ref. Piaggi et al. [33]) were performed to generate a series of configurations of the water-feldspar interfaces, covering all 13 K-feldspar terminations. The energies and forces for these configurations were then calculated using SCAN DFT to expand our dataset reported in Ref. Piaggi et al. [33]. The resulting dataset, which included new configurations and their corresponding energies and atomic forces, was used to train a set of four MLPs. Based on the newly trained MLPs, additional configurations were explored, and this cycle was repeated iteratively until a high-accuracy MLP with SCAN-level precision was obtained. Figures S13 and S14 show that our MLPs have excellent and uniform accuracy across all studied terminations. Settings of DeePMD-kit are identical to those used in Ref. Piaggi et al. [33].
MD simulations
We performed MD simulations and enhanced sampling simulations using LAMMPS[35, 48] interfaced with the PLUMED plugin[50, 47]. All simulations employed the custom-trained MLPs described above to model atomic interactions. We used the standard atomic weights for the masses of all elements (K, Al, Si, O) except H, for which we used a mass of 2 grams/mol, in order to improve the stability of the integration of the equations of motion. Simulation boxes with periodic boundary conditions in all directions were used. The time step for the integration of the equations of motion was 0.5 fs. The simulation models were first energy-minimized using the conjugate gradient (CG) method and then equilibrated using MD simulations in the ensemble. A stochastic velocity-rescaling thermostat[2] with a 0.1 ps relaxation time was used to control the temperature. For studies of water structure on multiple surfaces, simulations were performed at 290 K (supercooling =18 K) for more than 3 ns. During the simulations, the K, Al, and Si atoms were constrained to their initial positions by a harmonic spring with force constant of 20 . The trajectory was saved every 1000 steps.
As standard MD simulations cannot capture the nucleation process within affordable simulation times, we employed an enhanced sampling method, namely steered MD simulations guided by the Steinhardt order parameter[43, 51], as implemented in PLUMED[50, 47]. The oxygen atoms of water molecules were used to calculate the Steinhardt order parameter, which quantifies local structural ordering. A moving restraint with a force constant of was applied to this collective variable. The simulations were performed at 300 K (supercooling =8 K). The target value of the restraint was gradually increased from 0.05 to 0.2 over a 5 ns simulation. The order parameter was recorded every 1000 steps to monitor the progress of nucleation throughout the simulation, and the trajectory was saved every 10000 steps. After the steered MD simulations, configurations were extracted from the trajectory and used as the initial models for seeding simulations without bias. A very similar steered MD protocol was extensively validated for studying ice-binding surfaces in Ref. Naullage et al. [30]. The simulations reported in that work show that while using biased simulations may result in a different polymorph than in the unbiased simulations, this impacts mostly the ice that forms far from the surface.
The free energy profile of ice nucleation was calculated using the umbrella sampling method. The Steinhardt parameter was used as collective variable, and 41 windows in the range with a width of 0.0015 were used. In each window, the parameter was restrained by a harmonic potential with a spring constant of . Each simulation was performed for 3 ns. The time series of the parameter for all windows were then used to construct the free energy profile via the weighted histogram analysis method (WHAM), following the implementation of Grossfield[26, 37, 11].
Visualization and analysis of the MD trajectories were performed using VMD[18]. To identify ice-like atomic environments we employed the Identify Diamond Structure modifier in OVITO[44] using the oxygen atoms. We quantified the number of ice-like water molecules by counting the number of oxygen atoms classified as having a diamond-like structure, and we ignored their first and second neighbors with imperfect structures, which are sometimes included in this type of calculation. The results are fully consistent with the values obtained using the Polyhedral Template Matching modifier[27, 34].
Obtaining thermodynamic properties using classical nucleation theory
The effective surface free energy for the heterogeneous nucleation process can be computed from an equation that closely follows its counterpart for homogeneous nucleation, namely,
where is an effective surface free energy that combines information about the water-ice surface free energy and the contact angle, is the density of ice, and is the chemical potential difference between liquid water and ice. and are easily computed and were taken from previous SCAN-MLP-based simulations reported in ref. Piaggi et al. [34]. According to classical nucleation theory, is related to the nucleation free energy barrier, , as follows,
Another way to characterize the ice nucleation process is through the heterogeneous nucleation rate, which can be calculated as,
where molecules/nm2 is the density of liquid water at the first layer in contact with the substrate’s surface, is the Zeldovich factor, is the attachment rate of molecules to the critical nucleus, with being the Boltzmann constant and the temperature. and were taken from ref. Piaggi et al. [34].
Estimation of nucleation rates in experiments
We estimated the experimental ice nucleation rate using Fig. 2A of Ref. Kiselev et al. [23]. In this figure, the nucleation sites of individual ice crystals were plotted during eight subsequent nucleation–evaporation cycles on the defects of a feldspar (001) cleavage surface within 100 s. Approximately 200 sites were marked over the eight cycles in the observed region. The area of the observed (001) cleavage surface is . Clearly, the exposed (110) surface area is much smaller than that of the (001) surface, and thus the area of (110) surface was estimated using a scaling factor, , such that the effective exposed area of (110) surface is . Therefore, the ice nucleation rate, , can be estimated as
where represents the average number of nucleation sites per cycle, and is the duration of each nucleation–evaporation cycle. The estimated nucleation rate is . The scaling factor is much smaller than 1, and assuming to be in the order of to , we computed the range , which is highlighted in Fig. 3I.
Acknowledgements
We are grateful for the computational resources provided on the Hyperion cluster at the Donostia International Physics Center (DIPC) and on MareNostrum 5 ACC at the Barcelona Supercomputing Center, granted through the Spanish Supercomputing Network (RES) allocations FI-2024-2-0026 and FI-2024-3-0028. We thank Prof. Hu Qiu from Nanjing University of Aeronautics and Astronautics for valuable suggestions on the data presentation in this work. W.Z. was funded by the ’María de Maeztu’ Units of Excellence program No. CEX2020-001038-M. P.M.P. acknowledges funding from the Marie Sklodowska-Curie Cofund Programme of European Commission project H2020-MSCA-COFUND-2020-101034228-WOLFRAM2.
Author contributions:
W.Z. performed research. W.Z. and P.M.P. designed research, analyzed data, and wrote the manuscript. P.M.P conceived the project.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
The scripts used for generating slab surfaces, the dataset for training the MLPs, the input files for MLP training, the final trained MLP, as well as the input files for the MD simulations and DFT calculations reported in this work, are openly available on Zenodo (DOI https://doi.org/10.5281/zenodo.17610332)[61].